RETARDATION OF PLATEAU-RAYLEIGH INSTABILITY: A DISTINGUISHING CHARACTERISTIC AMONG PERFECTLY

WETTING FLUIDS

JOHN MCCUAN

Abstract. We consider a cylindrical film of fluid adhering to a rigid cylinder of fixed radius. The main result is to give the critical (maximum) length for which such a film of given thickness can be stable. The problem is considered both when the cylinder remains stationary and when the fluid and the cylinder are co-rotating at a fixed angular velocity.